OpenStudy (anonymous):

Use implicit differentiation to find dy/dx for y^2 = 2xy – x^2

OpenStudy (anonymous):

best way to explain it is to treat y like a variable x .... which gets you: 2 y y'=2xy'+y-2x

myininaya (myininaya):

2yy'=2xy'+2y-2x

myininaya (myininaya):

Solve the above for y' and u are done

OpenStudy (anonymous):

yea that forgot some brackets

OpenStudy (anonymous):

is the 2yy 2y^2?

myininaya (myininaya):

?I dont understand the question

OpenStudy (anonymous):

is the 2yy you wrote suposed to be 2y^2

myininaya (myininaya):

I think you are talking about 2yy' this does not equal 2yy

OpenStudy (anonymous):

okay I got 2y-2x

myininaya (myininaya):

do you know how to solve this equation for x: x+4=2x-3

OpenStudy (anonymous):

2yy'-2xy'=2y-2x y'(2y-2x)=2y-2x y'=1

myininaya (myininaya):

Eric is right